Find the volume of 5 molal H2 SO4 solution(d=1.2g / ml) which should be added to 100ml of 98% H2 SO4 solution (d=1.8 g/ ml) in order to obtain 10 molal H2SO4 solution.
 

Grade:11

1 Answers

Himanshu Rathour
5 years ago
Let v litre 5 molal H2SOis required
5 molal H2SO=      5moles of H2SO4
                                     1 kg of solvent
=5moles of H2SO4        {density = mass/volume
    1/1.2 litre of solvent
 
Molarity=.    No. Of moles
                         Volume
=6 molar solution
Now if we take v litre of this solution
Then, no. Of moles of H2SO4 in v litre=6v
For II solution
98% H2SO4 =98ml H2SO4 in 100 ml
Wt. Of H2SO4 =98*1.8
Moles =98*1.8/98=1.8
Now if we mix both solutions then total moles=1.8+6v
Total volume Of solvent=0.1+v
10=     6v+1.8      
            0.1+v
V=0.2litre